How many grams are in 0.2 moles of H2O


1
Expert's answer
2021-05-14T02:34:57-0400

How many grams are in 0.2 moles of H2O?

The number of moles which is n= 0.2mol

The molar mass of water which is m= 18g/mol

Therefore, the mass of water which is w=n*m=0.2*18

Grams of H2O in 0.2 moles= 3.6grams